### v1

Abstract
Object-Centric Process Mining (OCPM) enables business process analysis from multiple perspectives. For example, an educational path can be examined from the viewpoints of students, teachers, and groups. This analysis depends on Object-Centric Event Data (OCED), which captures relationships between events and object types, representing different perspectives. Unlike traditional process mining techniques, extracting OCED minimizes the need for repeated log extractions when shifting the analytical focus. However, recording these complex relationships increases the complexity of the log extraction process. To address this challenge, this paper proposes a method for extracting OCED based on PM, a well-established process mining framework. Our approach introduces a structured framework that guides data analysts and engineers in extracting OCED for process analysis. We validate this framework by applying it in a real-world educational setting, demonstrating its effectiveness in extracting an Object-Centric Event Log (OCEL), which serves as the standard format for recording OCED, from a learning management system and an administrative grading system.
Keywords:
Object-Centric Process Mining, Methodology, Log Extraction
1 Introduction
Object-Centric Process Mining (OCPM) is a data-driven approach that enables multi-perspective process analysis, allowing for a more comprehensive understanding of complex workflows [2, 17]. Unlike traditional process mining techniques that focus on a single perspective - such as analyzing the learning process from the student perspective alone - OCPM considers multiple object types simultaneously, revealing hidden dependencies and bottlenecks that might remain undetected when examining processes from a singular viewpoint [16]. For example, in an educational setting, the learning process can be analyzed not only from the perspective of students but also from that of teachers and course groups, enabling insights into instructor-student interactions, group-based collaboration, and administrative workflows.
To apply OCPM in practice, Object-Centric Event Data (OCED) [7] must be extracted to capture events in relation to multiple object types. In an educational setting, typical object types include students, teachers, and groups, each of which plays a distinct role in the learning process. Unlike traditional event logs, OCED reduces the need for repeated log extraction when shifting between analytical perspectives. Several formats exist for representing OCED, including the Object-Centric Event Log (OCEL) [2], Data-Aware Object-Centric Event Logs (DOCEL) [9], Event Knowledge Graph (EKG) [6], and Temporal Event Knowledge Graph (tEKG) [11], which can be transformed into one another [10, 11]. However, extracting OCED from operational systems remains inherently complex, as it requires capturing relationships between events and all possible object types while ensuring proper data integration.
Despite recent advancements, OCED extraction from information systems continues to pose significant challenges, necessitating new methods and guidelines to support researchers and practitioners. While prior studies [1, 3, 19] have addressed specific challenges in OCED extraction, existing solutions are often system- or platform-dependent, tailored to particular environments such as OnProm [19] or designed for extracting logs from Enterprise Resource Planning (ERP) systems [1, 3].
To bridge this gap, this study proposes a structured and system-agnostic framework for OCED extraction, addressing the complexities associated with capturing multi-perspective event data. Our approach extends PM [18], a well-established process mining framework, offering a systematic framework to support data analysts and engineers in extracting event logs for process analysis. To demonstrate its practical applicability, we applied our framework in the educational domain, extracting OCEL from a learning management system and an administrative grading system. This enables the simultaneous analysis of educational paths from both the group and student perspectives, allowing for a deeper understanding of learning dynamics and collaborative processes.
The remainder of this paper is structured as follows. Section 2 elaborates on previous process mining methodologies, provides an overview of two primary OCED formats, and summarizes existing object-centric event data extraction approaches. Section 3 presents our OCED extraction method. Section 4 details the application of our method for extracting OCEL from learning management systems. Finally, Section 5 concludes the paper.
2 Background
This section provides an overview of existing process mining methodologies and peer-reviewed approaches for object-centric event data (OCED) extraction. In addition, it explores two primary approaches for representing OCED.
2.1 Process Mining Methodologies
Process Diagnostics Method (PDM) [4], the L* Life-Cycle Model [5], and PM2 [18] are three well-known process mining methodologies followed by different people to apply process mining in practice.
The Process Diagnostics Method (PDM) [4] focuses on rapid, high-level analysis of business processes without requiring domain-specific knowledge. It examines control flow, performance, and organizational aspects based on event logs, providing quick insights. However, PDM lacks explicit iterative refinement of the analysis [18], and is less suitable for long-term process optimization efforts [4].
In contrast, the L* Life-Cycle Model [5] follows a structured, stepwise approach that includes planning and justification, event log extraction, process model discovery, integration, and operational support. It defines three distinct project approaches: goal-driven (focusing on achieving predefined objectives), question-driven (addressing specific business queries), and data-driven (exploring available data for insights). However, this methodology is effective solely for structured processes where a single integrated process model can be developed.
To overcome the limitations of both PDM and L*, PM2 [18] introduces an iterative process mining methodology that incorporates stakeholder involvement and a structured roadmap for continuous process improvement. Unlike L*, which primarily targets structured processes, PM2 accommodates both structured and unstructured processes, allowing for evolving business-related questions and iterative refinements.
The PM2 methodology consists of six key stages [18]: (i) Planning, which involves defining project goals, such as performance improvement or compliance verification, formulating research questions (we refer to as business-related questions throughout the paper), and identifying relevant business processes. This stage also involves assembling a multidisciplinary project team, including business experts, system specialists, process analysts, and business owners; (ii) Extraction, where event data is collected from information systems, considering data quality, scope, and process knowledge transfer; (iii) Data Processing, which prepares event logs by defining case notions, filtering, aggregating, and enriching event data; (iv) Mining and Analysis, where process discovery, conformance checking, and performance enhancement techniques are applied; (v) Evaluation, which includes interpreting analysis results, validating insights with domain experts, and refining business-related questions; and (vi) Process Improvement and Support, where process changes are implemented, or operational support is provided based on the findings.
These methodologies offer structured approaches for applying process mining in practice and are well-suited for extracting traditional event logs from source systems. However, extracting OCED presents additional complexities due to the involvement of multiple object types, each requiring the capture of different events. As a result, analysts can easily overlook some necessary relationships, leading to repeated data extraction. This redundancy creates bottlenecks in the log extraction process. To address this challenge, we extend PM2 to enhance the efficiency of OCED extraction. Before introducing our approach in the next section, the following subsections provide an overview of OCED formats and discuss existing methodologies for OCED extraction.
2.2 Brief Overview of OCED
Object-Centric Event Data (OCED) provides a richer, more realistic representation of process event data by capturing multiple objects, their relationships, and interactions. Unlike traditional event logs, which focus on a single case notion (e.g., an order or a customer), OCED captures the relationships between different entities related to events. Figure 1 presents two metamodels for storing OCED: Object-Centric Event Logs (OCEL 2.0) [2] and Event Knowledge Graphs (EKG) [6].
(a) OCEL 2.0 metamodel
(b) EKG metamodel
Figure 1: Two sample metamodels for storing OCED
The OCEL 2.0 metamodel, shown on the left side of the figure in UML notation, defines an event as an atomic action that occurs in a business process, representing the execution of an event type (a.k.a, an activity), such as “place order” in Order-to-cash process. Each event type can have multiple event attributes, and each event can store different event attribute values. OCEL records objects as separate entities, where each object belongs to an object type. For example, the “order” can be an object type with multiple instances representing different orders. Each object type can have multiple object attributes, and each object can store different values for these attributes at specific points in time. Objects can also be related to other objects through object-to-object relations; for example, an order can contain multiple items. As illustrated in the metamodel, events can be associated with multiple objects and vice versa through event-to-object relations.
EKG, on the right side of the figure, provides an alternative approach for capturing OCED by representing events, entities, logs, and classes as nodes with specific relationships among them. Events and classes in EKG correspond to events and event types in OCEL, respectively. Entities in EKG represent objects, with object types stored as a property of the entity node called EntityType. Object-to-object relations are captured as edges between entity nodes labeled REL, while event-to-object relations are represented as edges between event and entity nodes labeled CORR.
There are more variations capturing OCED data like temporal EKG [11], which improves EKGs by tracking how object attributes change over time so events can be analyzed using the correct object information at each point in time, and Data-aware Object-Centric Event Logs (DOCEL) [9], which extends OCEL by tracking attribute changes over time, linking attributes to objects and events, and allowing attributes to have multiple values.
2.3 OCED Extraction Initiatives
Several approaches have been developed to extract OCED from information systems, particularly Enterprise Resource Planning (ERP) systems. Here, we exclude methods that focus solely on transforming OCED formats from other formats or traditional event logs since this study focuses on proposing a method for OCED extraction directly from information systems.
Berti et al. proposed a method for extracting event data from SAP ERP systems using OCEL [3]. Their approach begins by constructing a Graph of Relations (GoR) to map relevant business process tables. Once these tables are mapped, data are extracted and stored in the OCEL format. To streamline this process, they developed a Python-based tool that identifies relevant tables and their relationships. The extracted data can subsequently be used for OCPM or converted into traditional event logs.
Building on this approach, Berti et al. conducted a case study focusing on the purchase-to-pay (P2P) process [1]. Similar to their previous work [3], this study employed graph-based techniques to model relationships within SAP’s complex data structures. The methodology utilized an object interaction graph to visualize object relationships and applied the PM2 framework to structure the extraction and analysis process within a well-defined methodological context.
Beyond SAP-specific solutions, Xiong et al. introduced a method for extracting OCEL from relational databases using a Virtual Knowledge Graph (VKG) approach [19]. Their method extended the OnProm framework to support both OCEL and the traditional XES format. By leveraging ontology-based data access (OBDA) and the VKG system Ontop, they defined domain ontologies and mappings, enabling event log extraction via SPARQL queries. Their approach was validated using the Dolibarr ERP system, demonstrating how relational data can be transformed into structured OCEL logs.
Unlike these system-specific approaches, this paper presents a generalized framework for OCED extraction that is adaptable across various information sources. Our methodology ensures broader applicability and ease of implementation across different systems by providing a step-by-step process that does not rely on platform-dependent techniques.
3 The OCPM2 Methodology
This section introduces OCPM2, an Object-Centric Process Mining methodology defined as an extension of PM2 [18] to facilitate the extraction of Object-Centric Event Data (OCED) [7]. We define our method by extending PM2 to ensure the reuse of a well-established and previously tested methodology for event log extraction, originally designed for traditional process mining. Our approach retains the clarity and efficiency of PM2 while making it suitable for extracting object-centric event data from diverse information systems.
The methodology consists of twelve stages, where the related stages are grouped into phases, as indicated by dashed rectangles in Figure 2. The process begins with the Planning stage, followed by the Domain Modeling, Log Extraction, and Analysis Iteration phases, and concludes with the Process Improvement & Support stage. Each stage produces specific artifacts, as outlined in the figure. The following sections elaborate on these stages, the generated artifacts, and the defined phases. The filled-colored phases and stages are identical to those in the PM2 methodology briefly described in Section 2 (we refer readers to the PM2 [18] for further details.)
Figure 2: An overview of the OCPM2 methodology extending the PM2 methodology
3.1 Planning
The Planning stage follows PM2, producing the following key artifacts: i) prioritized business-related questions (denoted by ? in the Figure 2), ii) a list of information systems ( IS) from which data will be extracted to answer these questions. The minor difference from PM2 is that the questions are prioritized, allowing for prioritizing the object type extraction later.
3.2 Domain Modeling
The Domain Modeling phase consists of five stages (stages 2 to 6 in Figure 2), identifying and conceptualizing the information needed for the extraction of OCED.
The phase begins with Object Type Identification to determine all object types needed to address the questions elicited in the Planing. Because an object type may be required for multiple questions, we propose documenting this using a Question-to-Object Type Matrix (Q2OT). This matrix represents questions as rows and object types as columns, marking intersection cells to indicate the required object types for each question.
Next, in the Conceptual Modeling stage, process analysts and business experts collaborate to define a model ( CM) that captures the relationships between object types. Business experts provide domain knowledge, while process analysts structure this information using a conceptual modeling language to represent key relationships. Among these, recognizing the “is-a” relationship enables capturing OLAP operations such as drill-down and roll-up in OCPM based on OCEL [13, 12].
The Activity Identification stage follows, eliciting different activities (or event types) to which the identified object types and questions are related.
The Extraction Matrix Definition stage defines the extraction matrix (X), where the relationships between object types and activities are documented. When an “is-a” relationship exists among object types, separate columns can specify whether an activity relates to the supertype class or a subtype class. Various documentation approaches can be adapted based on specific perspectives. From a system perspective, CRUD (Create, Read, Update, and Delete) operations are suitable for specifying the type of relationship between an activity and an object type. However, from a process perspective, CRUD may not fully capture the dynamic nature of object interactions. Instead, documenting the quantity range of objects for each object type provides a more process-oriented approach. This method captures the expected number of objects per event and supports data validation by ensuring that the extracted log adheres to these expectations. The details of these adaptations can be delegated to design and development teams.
Finally, in the Extraction Prioritization stage, business, process, and system experts collaborate to prioritize the extraction of objects and events based on their i) feasibility, ii) data availability, or iii) importance. Given the potentially large number of object types and activities, extracting all data at once may not be feasible due to data being fragmented across multiple sources and the need to check data quality step-by-step to catch errors early. To manage this, we propose a prioritization artifact as the output of this stage, including a set of prioritized objects and activities ( Pr). Note that new object types can be identified during this phase, requiring the phase to be repeated to have a more holistic understanding of the domain.
3.3 Log Extraction
The Log Extraction phase consists of three stages (Stages 7 to 9 in Figure 2) as follows:
In the Implementation stage, developers build pipelines to read data from source systems, transform them into the desired OCED format, and load them into destination storage. If OCEL is the target format, objects should be processed first based on their list in the extraction matrix. When an “is-a” relationship exists among object types, we propose storing objects at the supertype level with an attribute distinguishing subtypes. This structure follows the single-table inheritance technique used in Object-Relational Mapping (ORM) [8]. This ensures referential integrity and facilitates efficient querying and aggregation, allowing drill-down operations [12]. Subsequently, developers implement extraction logic for object-to-object relationships, determined by the conceptual model (stage 3), in parallel with extraction logic for events based on the extraction matrix. Once event extraction is complete, they implement extraction logic for event-to-object relationships. Please note that similar mapping can be done for EKG [6] and tEKG [10] as the OCED formats are transformable to each other.
The Extraction stage is separated from Implementation because it enables the development of data pipelines to extract, clean, and integrate data for process mining as well as testing the result before final extraction in a production environment.
In the Verification stage, correctness can automatically be checked by deriving a matrix from the extracted event data. This matrix shows the number of object types for each event type, ensuring that documented relationships in the Extraction Matrix are properly implemented.
Although the log extraction process has been adapted to generate OCED, the following stages remain identical to PM2 [18] since the subsequent stages of the process mining apply regardless of the log format. Additionally, the extracted OCED can be flattened to enable the use of traditional process discovery techniques.
3.4 Analysis Iteration
The Analysis Iteration phase in PM2 supports gaining insights from extracted event data through two stages: Mining & Analysis and Evaluation. The Mining & Analysis stage includes process model discovery ( PM) from the input log ( L), conformance checking, and process enhancement based on analytical insights ( A). This can yield compliance findings ( C) and performance findings ( P). The Evaluation stage results in process improvement ideas ( I) or refined/new process-related questions ( ?+), which may prompt further analysis or new data extraction.
3.5 Process Improvement & Support
The Process Improvement & Support stage applies the insights from process analysis to optimize workflows and enhance performance. It involves defining strategies, supporting implementation, and monitoring changes.
4 Case Study
This section presents the application of the OCPM methodology in executing a process mining project at the Department of Computer and Systems Sciences, Stockholm University. We elaborate on each stage and phase below.
4.1 Planning
This project was initiated to explore how process mining can provide teachers with data-driven insights crucial for improving their educational content and tasks’ efficiency. To determine the most important questions and their relative significance, we conducted interviews with 10 teachers. The interviews were recorded, transcribed, anonymized, and analyzed to extract key questions and assess their perceived importance. The outcome of this stage consisted of two artifacts: a set of prioritized questions compiled from the interviews ( ?) and a list of relevant information systems ( IS). The four most important questions identified were:
  • Q1: What learning paths do students typically follow when accessing educational materials, such as files, pages, and folders, throughout a course?
  • Q2: How do students submit (or resubmit) individual and group assignments during the course?
  • Q3: Do students who usually take the lead in submitting group assignments on behalf of the group tend to achieve higher final grades (exam and assignment grades)?
  • Q4: How is students’ exam success related to their frequency of accessing course materials, such as files, pages, and folders?
Students often follow different routes when accessing educational materials, influencing their engagement and performance. Thus, analyzing students’ learning paths (Q1) provides insights into how these behaviors influence the outcomes while also identifying the challenging areas that affect their success [15]. Examining assignment submission patterns (Q2) helps to assess student engagement and identify potential challenges in meeting deadlines. Furthermore, analyzing whether high-achieving students demonstrate different behaviors, such as taking the lead in submitting group assignments (Q3), provides insights into collaborative learning strategies, such as encouraging role rotation. Finally, analyzing the relationship between course material access and exam performance (Q4) can help educators identify patterns that distinguish successful students from those who struggle.
Although we identified numerous questions to guide the project, we present only a subset here to illustrate the application of the methodology. The information systems identified at this stage included 1) the learning management system, Moodle [14], and 2) the website that records students’ grades.
4.2 Domain Modeling
Object Type Identification: From the questions above, we identified the following object types: Student, File, Page, Folder, Assignment, Group, Course, and Exam. These object types were identified by extracting key nouns from the questions that belong to the domain vocabulary. These object types model the domain of interest. Table 1 presents the Question-to-Object Type matrix (Q2OT), derived from the previous stage.
Student File Page Folder Assignment Group Course Exam
Q1 * * * * *
Q2 * * * *
Q3 * * * * *
Q4 * * * * * *
Table 1: Question-to-Object Type matrix supporting object type identification
Conceptual Modeling: Figure 3 presents the conceptual model using UML notation, illustrating the relationships between object types. This model includes additional object types beyond those represented in the Q2OT matrix, reflecting the iterative nature of our approach. For example, the object type Teacher was later identified by reviewing the Moodle documentation as a relevant entity within the process. Additionally, User was introduced as a supertype encompassing both Teacher and Student as explained in Section 2.
Figure 3: The conceptual model illustrating object type relationships
Design choices also influence object type identification. For instance, Grade was modeled as an attribute of both Submission and Exam Sheet, though an alternative approach could define it as a separate class. The model also includes grayed-out classes, such as Exam Sheet, Submission, Module, Activity Module, and Resource Module. While some of these could be considered object types, their selection depends on modeling design choices. For example, Exam and Assignment are tightly related to Exam Sheet and Submission, respectively, due to the decomposition relation nature. Thus, we have not considered defining separate object types for Exam Sheet and Submission, which are parts of Exam and Assignment.
Activity Identification: Activities were directly identifiable from the questions by examining the leading tasks (verbs) in the questions. For example, view file, view page, and view folder were relevant for Q1, while submit assignment and resubmit assignment addressed Q2. Similarly, set assignment grade, set exam grade, and submit assignment were relevant for Q3. In addition, set exam grade, view file, view folder, and view page were relevant for Q4.
Extraction Matrix Definition: Table 2 documents relationships between identified activities and object types. In this version, we recorded the possible number of instances for each object type within the table’s cells. To provide more detailed documentation, we included User, the supertype of Teacher and Student, as a separate column.
User Exam File Page Folder Assignment Group Course
User Teacher Student
view file 1 1 1
view page 1 1 1
view folder 1 1 1
submit assignment 1 0..* 1 0..1 1
resubmit assignment 1 0..* 1 0..1 1
set assignment grade 1 1 0..* 1 0..1 1
set exam grade 1 1..* 1 1
Table 2: The extraction matrix
Extraction Prioritization: In our study, Assignment was prioritized highest, followed by File, based on teachers’ feedback, as it was the primary means of assessing student progress throughout the course before the final exam, and File was mainly used for uploading educational materials to support learning. Extraction was planned accordingly in the following sequence: submit assignment, resubmit assignment, set assignment grade, view file, view page, view folder, and set exam grade. Ultimately, we implemented all the extractions so their order does not affect the final result.
4.3 Log Extraction
Implementation: We set up development and test environments to build and validate pipelines for extracting data from the systems, integrating them based on the extraction matrix, and loading it into an OCEL format. When loading data into OCEL, we defined User as a single object type, with Teacher and Student modeled as subclasses. This approach facilitated the discovery of process models at a general level while enabling drill-down analyses to distinguish teacher- and student-specific process behaviors.
Extraction: Data extraction was performed on a single course over one year. This timeframe provided a comprehensive dataset of interactions between students, teachers, and learning materials while ensuring manageability. By focusing on one course, we captured the full lifecycle of learning activities, including assignments, submissions, grading, and resource interactions.
Verification: We developed a script to systematically verify the accuracy of the extracted event data by mapping activities to object types. Figure 4 presents this structured validation process. In this figure, the expected relationship between Group and set assignment grade is missing. This discrepancy arises because Moodle logs a separate set assignment grade event for each student, even when a teacher grades a group assignment collectively. Automated verification helped identify such inconsistencies, highlighting areas for improvement in future log extractions by developers. Furthermore, the relationship between Teacher and Student is not directly visible because the types of objects were recorded at the User level. However, drill-down analysis can reveal these relationships as needed.
Figure 4: The verification matrix
4.4 Analysis Iteration
We analyzed the extracted logs to answer the identified questions. The results were evaluated by the course responsible to ensure accuracy.
For Q1, Figure 5 illustrates an Object-Centric Directly-Follows Graph (OC-DFG), discovered by filtering the extracted log, drilling down page objects into individual page names, and unfolding view page events for each page. This graph visualizes the sequential flow of actions during the course from the perspectives of the User and Course object types.
Refer to caption Refer to caption
Figure 5: Object-Centric Directly-Follows Graph (OC-DFG) illustrating how students visited some pages
In this process, the page names for given codes are as follow: A1P1: History of BPM, A1P2: Business Process Models, A1P3: Business Process Enactment, A1P4: Syntax and Semantics, A1P5: Control-flow Patterns, A1P6: Business Process Complexity, and L1P1: Introduction to Process Tree. The model reveals that students sequentially accessed five pages: A1P1 to A1P5. Additionally, a loop between A1P5, A1P6, and L1P1 suggests students frequently revisit these pages, indicating potential room for improvement.
As an example of an analysis related to Q2, we refer to Figure 6, which presents an OC-DFG illustrating how individual and group assignments were submitted. The OC-DFG was discovered by filtering the log based on submission activities and events, drilling down user objects into user types (separating teachers and students) and assignments into their specific names (distinguishing different assignment object types). Additionally, the log was unfolded to include related events for these objects.
Refer to caption Refer to caption
Figure 6: OC-DFG illustrating the overall submission process for individual and group assignments in addition to a zoomed part
The upper part of the figure displays the entire process, and the lower part presents a zoomed-in version that crops the beginning of the process to enhance readability for further elaboration. The upper figure may not be easily readable (which is intentional, as we do not elaborate on details), but it effectively illustrates the sequential flow within the process. This sequence is influenced by the structured delivery of educational materials and the course setup.
In the lower part of the figure, we observe that 23 students submitted the assignment, as it was a group submission, where both the group and individual students were linked to the event. However, we also see that the grading process involved 134 students. This discrepancy highlights the issue discussed in the verification section regarding the lack of a direct relationship between group and set assignment grade.
Although, due to space constraints, we cannot present all analysis results in detail, extracting OCELs enabled us to investigate all the questions raised while interviewing teachers at the beginning of the project.
4.5 Process Improvement & Support
Our analysis revealed that students frequently revisited pages related to control-flow perspectives, process complexity, and process trees, indicating possible challenges in comprehension or the sequence in which these topics are introduced. To address this, we propose shifting the process tree discussion to follow control-flow perspectives and introducing process complexity at the end. This adjustment may reduce unnecessary back-and-forth navigation, making the learning process more structured. Further investigation is needed to understand the underlying reasons for these recurring visits, providing insights that could help refine the course.
5 Conclusion
This paper introduced OCPM2, an extension of the PM2 methodology, designed to support Object-Centric Process Mining (OCPM) by providing a structured approach for Object-Centric Event Log (OCEL) extraction. Our methodology enhances existing process mining frameworks by addressing challenges associated with multi-object process modeling, ensuring compatibility with structured process analysis techniques.
By incorporating domain modeling, OCED extraction, and iterative process analysis phases, OCPM2 provides a system-agnostic framework for object-centric process mining. The methodology reduces the complexity of extracting and analyzing multi-object event data, minimizing redundant log extraction efforts required for case-centric process mining.
To validate our approach, we applied OCPM2 in an educational setting, extracting and analyzing OCELs from a learning management system and an administrative grading system. Our results demonstrated the practical usefulness of OCPM2, i.e., its ability to effectively reveal student navigation patterns and assignment submission processes. The case study also exposed limitations in existing system logs, reinforcing the need for improved object-to-activity relationships in event data.
The case study exposed limitations in system logs, particularly in capturing accurate object-to-activity relationships. A key issue was the missing link between Group and set assignment grade, causing Moodle to log separate grading events per student, even for collectively graded assignments. This discrepancy resulted in a mismatch between the number of students who submitted the assignment and those involved in the grading process. Automated verification identified these inconsistencies, helping developers to extract correct event logs in future extractions.
Future research will focus on enhancing automated verification mechanisms, refining multi-dimensional process analysis operations, and expanding OCPM2’s applicability across other domains, such as healthcare and financial services. OCPM2 contributes to the advancement of Object-Centric Process Mining by providing an orchestrated approach to the mining task, with the potential for further enhancements such as automation and reuse.